OpenStudy (anonymous):

How would you solve: -2≤x-1≤0

OpenStudy (akshay_budhkar):

add 1 throughout the equation

OpenStudy (anonymous):

\[-1\le x \le 1\]

OpenStudy (akshay_budhkar):

\[-2+1\le x-1+1\le0+1\]

OpenStudy (akshay_budhkar):

and your answer will be as jakeb gave :)